5) Melanie read an advertisement from the Cell Phone Giants (CPG, for short, and yes, we're using a fictitious company name) that she thinks is too good to be true. The CPG ad states that customers in Los Angeles get average data download speeds of 4 Mbps. With this speed, the ad states that it takes, on average, only 12 seconds to download a typical 3=minute song from iTunes. Melanie thinks that it takes longer the time to download a typical 3=minutes song from iTunes; therefore she decides to gather evidence. Melanie collects a random sample of 45 downloads by using her friend’s phone to download her song from iTunes according to the randomly selected days, times, and locations. Melanie’s sample of size 45 downloads has an average download time of 13.5 seconds. The standard deviation for the sample is 3.2 seconds. What evidence did she finds?

Find the standard Error

Find the test statistics.
Answer #1

Answer)

Null hypothesis Ho : u = 12

Alternate hypothesis Ha : u > 12

Standard error = standard deviation/√n = 3.2/√45

Standard error = 0.47702783519

Test statistics t = (sample mean - claimed mean)/(standard error)

t = (13.5 - 12)/0.47702783519 = 3.144